Draw a histogram to represent the probability distribution for all the sums of TWO five-sided dice. (Hint: the sides of the dice are 1,2,3,4 & 5)
Answer #1

here total number of outcomes =5*5 =25 (as 5 outcome on one dice)

P(X=2) =1/25 (as one outcome (1,1) when sum is 2)

P(X=3)=2/25 (two outcomes (1,2),(2,1) are there)

P(X=4)=3/25

P(X=5)=4/25

P(X=6)=5/25

P(X=7)=4/25

P(X=8)=3/25

P(X=9)=2/25

P(X=10)=1/25

below is the histogram:

P(x) 1/4 1/5 3/20 1/10 1/20 10 P(x)